2019, 200 KGs of oranges are produced at Rs.40/kg, 500 KGs of apples are produced at 120/kg, and 200 Glasses of orange juice are produced at Rs.150/glass.

In 2020, 220 KGs of oranges are produced at Rs.50/kg, 450 KGs of apples are produced at 140/kg, and 230 glasses of orange juice are produced at Rs.135/glass.

 

Assuming 2019 is the base year; calculate the real GDP in 2019 



1
Expert's answer
2021-05-04T18:43:04-0400

It is note that in the base year, real GDP is equal to nominal GDP


So, we have to first calculate nominal GDP of 2019 assuming 2019 is the base year;


Nominal GDP = Sum of Price"\times" Quantity of each good


"= 200\times 40 + 500\times 120 + 200\times 150"

"= 8000 + 60000 + 30000"

= $98,000


As 2019 is the base year, so real GDP = nominal GDP = $98,000